# DialSim: A Real-Time Simulator for Evaluating Long-Term Multi-Party Dialogue Understanding of Conversational Agents

### v4

@@ Appendix F Prompt for Response Generation
The prompt given to the conversational agent to answer questions using dialogue history is shown in Table 9. An example where the placeholders from Table 9 are filled with actual values can be found in Table 10.
@@ G.1 Time Limit
In DialSim, the time limit is a controllable parameter, giving developers the flexibility to conduct experiments with any chosen time constraint, or even without one. When a time limit is set, the experimental environment can impact performance. Consequently, depending on the environment in which the conversational agent is deployed, this could serve as a criterion for selecting the agent with relatively better performance. It is important to note that the primary objective of DialSim is not to evaluate the inference speed of LLMs, but rather to assess the end-to-end performance of conversational agents, where techniques like model sharding and tensor parallelism can be a part of the conversational agent to decrease the response latency if needed.
@@ G.1.2 Average Time Interval Between Utterances
To account for this, we set a 6-second window as the response time limit for agents and conducted experiments to determine whether current models could meet this criterion. It is important to emphasize that the primary goal of these experiments was not to evaluate the absolute performance of the models but to showcase the range of analyses possible under time limits.
@@ G.1.3 Internet Speed
The performance of API-based models can be affected by internet speed. To analyze this, we conducted a comparative analysis of the response times between API-based models and open-source models. In our analysis of agents using OpenAI Embedding-Session Sum., we found that the API-based agents achieved average response times of 1.50 seconds for ChatGPT-4o-mini, 1.73 seconds for ChatGPT-3.5 and 2.69 seconds for Gemini 1.0 pro. In comparison, agents using open-source models showed average response times ranging from 2.06 seconds (Gemma 2B) to 7.15 seconds (Tulu2 70B). These results suggest that, even when accounting for both internet communication and model inference, remote API-based models are generally faster than open-source alternatives. This indicates that internet latency has a minimal impact on our evaluation.
@@ G.2 Question Format
LongDialQA is a dataset that includes pairs of questions, answers, and choices. The questions are available in three formats: template-based multiple-choice, natural language multiple-choice, and open-ended. Users can choose any of these formats to evaluate the agent’s performance.
@@
Additionally, we offer the option to ask questions in an open-ended format (e.g., “Who was going out with Paul in September 1994?”) without providing answer choices. This approach allows us to evaluate the agent’s ability to generate open-ended responses. The open-ended format is particularly useful for fan quiz-based questions, where some answers may require longer responses (e.g., Question: “Why did Monica and Chandler say they were late getting to the hospital?” Correct answer: “Monica went back for her jacket”).
@@ G.3 Number of Retrieved Dialogue History
By default, agents retrieved up to 20 utterances, 10 entire sessions, and 15 session summaries, depending on the storing method, though some LLMs with shorter context lengths retrieved fewer histories accordingly.
@@ Appendix I Experimental Results in an Unlimited Time Setting
The experimental results for different time limits are shown in Figure 6 and Figure 7. Figure 6 illustrates the performance over different time limits in the BM25-Session Entire setting, while Figure 7 displays the performance in the Oracle setting. Due to the high costs, time-based experiments with ChatGPT-4o, Gemini-1.5 Pro, and Claude-3 Opus were conducted exclusively in the Oracle setting. One key observation from the results is the performance of ChatGPT-3.5, ChatGPT-4o-mini, and ChatGPT-4o. These models demonstrated consistent performance with quick inference times, handling up to a 3-second limit in the BM25-Session Entire setting and up to a 1-second limit in the Oracle setting. Consequently, these models are optimal for tasks requiring real-time communication without delays.
@@
Figure 4: The result of asking ChatGPT-4o to explain Season 2, Episode 7 of Friends.
Figure 5: The result of asking ChatGPT-4o to explain Season 3, Episode 14 of Friends.
@@
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API ChatGPT-4o-mini 22.68 (2.12) 19.77 (2.02) 36.63 (1.82) 30.10 (1.44) 29.54 (0.71) 32.34 (0.58) 35.72 (0.81) ChatGPT-3.5 32.49 (1.72) 25.32 (1.20) 35.59 (2.12) 33.86 (1.09) 27.81 (0.40) 32.97 (0.86) 37.02 (1.13) Gemini 1.0 pro 3.49 (0.69) 25.87 (1.23) 30.72 (0.18) 38.16 (1.25) 37.42 (0.68) 32.09 (0.44) 36.30 (0.32) Open Tülu 2-70B 0.62 (0.13) 21.08 (0.70) 18.95 (1.07) 22.36 (0.65) 34.64 (0.69) 9.08 (1.00) 20.22 (1.48) Tülu 2-7B 0.53 (0.18) 15.58 (1.34) 22.26 (0.53) 29.99 (0.57) 16.84 (2.13) 21.48 (0.77) 28.69 (1.15) Llama3.1-70B 0.25 (0.07) 21.55 (0.93) 0.15 (0.12) 1.26 (0.31) 34.21 (1.59) 3.89 (1.05) 14.89 (1.74) Llama3.1-8B 21.30 (1.68) 12.80 (1.06) 25.50 (0.16) 18.56 (0.99) 23.10 (2.69) 25.48 (3.65) 20.75 (1.58) Mixtral-8x7B 1.95 (0.34) 15.91 (0.71) 34.52 (1.12) 16.83 (1.60) 17.45 (0.49) 34.98 (0.99) 13.83 (2.18) Mistral-7B 3.11 (0.21) 24.69 (1.82) 34.26 (0.60) 32.17 (1.39) 30.23 (0.62) 33.36 (0.56) 29.19 (1.54) Gemma-7B 16.40 (0.74) 21.40 (2.33) 19.74 (2.45) 16.67 (0.40) 24.50 (1.87) 20.22 (1.39) 16.12 (0.52) Gemma-2B 1.56 (0.06) 28.94 (0.35) 26.12 (2.22) 33.47 (1.41) 27.92 (0.68) 29.40 (1.79) 34.86 (3.20)
$\dagger$: Both ChatGPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
@@
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API ChatGPT-4o-mini 28.48 (1.01) 29.44 (0.62) 43.16 (1.37) 35.92 (2.50) 37.81 (0.30) 40.91 (0.37) 42.83 (1.12) ChatGPT-3.5 36.54 (0.32) 36.63 (0.57) 45.33 (1.00) 40.93 (0.13) 42.49 (1.24) 43.04 (0.82) 45.18 (0.56) Gemini 1.0 pro 2.42 (0.18) 35.11 (0.50) 48.90 (1.57) 40.91 (0.75) 44.72 (0.19) 46.63 (0.89) 45.82 (0.97) Open Tülu 2-70B 0.46 (0.09) 22.33 (1.00) 35.52 (0.89) 23.49 (1.16) 38.61 (1.02) 43.49 (1.27) 23.54 (0.52) Tülu 2-7B 0.32 (0.04) 25.86 (0.54) 27.95 (1.03) 36.60 (2.11) 22.13 (0.33) 29.50 (0.56) 35.51 (1.40) Llama3.1-70B 0.19 (0.07) 29.21 (0.56) 13.31 (0.94) 21.32 (5.22) 47.41 (0.93) 47.07 (1.32) 19.46 (1.73) Llama3.1-8B 21.87 (0.60) 22.03 (0.32) 37.94 (1.28) 29.16 (1.80) 27.76 (3.52) 37.67 (1.70) 26.67 (0.83) Mixtral-8x7B 1.53 (0.41) 19.63 (0.79) 34.35 (1.19) 16.07 (0.56) 20.02 (0.44) 30.44 (1.69) 12.43 (1.04) Mistral-7B 2.55 (0.09) 30.65 (0.45) 41.16 (1.26) 35.67 (1.68) 36.92 (2.13) 42.71 (1.24) 37.65 (2.42) Gemma-7B 17.81 (0.86) 21.58 (0.61) 25.62 (0.02) 12.20 (0.57) 24.88 (0.93) 24.38 (0.52) 15.70 (0.43) Gemma-2B 0.83 (0.16) 29.71 (0.69) 28.11 (1.14) 34.63 (0.94) 31.54 (0.65) 30.31 (0.16) 33.37 (0.27)
$\dagger$: Both ChatGPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
@@
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API ChatGPT-4o-mini 38.91 (0.99) 34.44 (0.52) 49.21 (0.12) 42.23 (1.57) 38.91 (0.74) 43.64 (0.42) 42.40 (0.99) ChatGPT-3.5 31.81 (1.33) 26.91 (2.30) 39.45 (1.40) 32.77 (1.31) 32.41 (0.96) 35.78 (0.74) 35.98 (1.75) Gemini 1.0 pro 28.36 (0.97) 28.10 (1.08) 39.90 (1.08) 34.11 (1.64) 34.26 (2.91) 30.93 (2.17) 33.96 (2.11) Open Tülu 2-70B 3.31 (0.32) 29.87 (0.65) 35.87 (2.56) 34.72 (1.63) 37.07 (0.72) 33.63 (1.32) 38.62 (1.94) Tülu 2-7B 1.57 (0.12) 28.93 (2.81) 28.72 (1.80) 30.86 (2.29) 34.55 (0.47) 31.04 (0.96) 32.12 (0.75) Llama3.1-70B 36.36 (0.68) 31.84 (1.29) 43.17 (0.99) 43.81 (0.94) 39.85 (2.08) 43.17 (0.68) 48.49 (0.97) Llama3.1-8B 28.78 (0.34) 29.89 (1.56) 34.70 (1.75) 33.93 (1.76) 31.63 (2.17) 32.91 (0.51) 35.59 (1.09) Mixtral-8x7B 42.19 (1.76) 31.84 (0.78) 46.47 (1.75) 32.31 (1.09) 35.51 (0.19) 41.24 (2.90) 34.18 (0.96) Mistral-7B 32.93 (0.59) 28.20 (1.17) 35.09 (1.76) 30.16 (1.82) 30.12 (1.45) 31.00 (1.93) 30.80 (1.75) Gemma-7B 18.78 (0.87) 22.26 (1.52) 23.62 (2.09) 19.83 (1.74) 25.07 (0.49) 22.48 (0.25) 20.08 (0.76) Gemma-2B 1.16 (0.26) 25.03 (1.54) 24.64 (1.31) 24.84 (2.05) 28.06 (1.38) 24.56 (2.60) 28.28 (1.94)
$\dagger$: Both ChatGPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
@@
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. Open Tülu 2-70B 0.31 (0.13) 17.03 (0.94) 15.20 (0.87) 18.45 (1.04) 26.89 (0.54) 6.92 (0.47) 12.86 (1.23) Tülu 2-7B 0.73 (0.29) 12.50 (1.73) 17.58 (1.14) 24.21 (1.09) 10.20 (0.21) 14.26 (0.92) 21.03 (0.58) Llama3.1-70B 0.51 (0.00) 27.84 (1.89) 0.60 (0.06) 13.84 (2.31) 35.67 (1.89) 1.53 (028) 20.90 (0.37) Llama3.1-8B 25.84 (1.16) 25.24 (0.30) 28.86 (1.01) 24.56 (0.99) 28.86 (1.10) 32.35 (1.51) 24.05 (1.36) Mixtral-8x7B 1.77 (0.19) 14.03 (0.12) 21.11 (1.07) 15.50 (0.68) 13.14 (0.83) 18.03 (0.44) 18.47 (0.55) Mistral-7B 2.34 (0.17) 22.29 (1.43) 27.08 (0.99) 24.15 (1.76) 25.17 (1.74) 26.76 (2.64) 23.81 (2.53) Gemma-7B 18.87 (1.43) 22.85 (0.81) 22.96 (1.34) 17.95 (0.62) 25.46 (2.08) 21.53 (1.00) 17.66 (1.31) Gemma-2B 0.78 (0.22) 22.99 (0.66) 25.48 (1.54) 25.86 (2.48) 25.08 (1.34) 25.21 (0.22) 26.14 (1.71)
$\dagger$: Llama3.1 supports up to 128k tokens, but we limited it to 8k tokens due to GPU VRAM limits.
@@
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. Open Tülu 2-70B 2.54 (0.21) 26.47 (1.91) 31.75 (1.71) 30.94 (2.41) 31.90 (1.25) 29.83 (1.03) 31.86 (2.13) Tülu 2-7B 1.15 (0.06) 28.20 (1.63) 27.64 (2.37) 27.78 (1.32) 28.98 (0.96) 25.03 (0.94) 29.08 (2.47) Llama3.1-70B 31.38 (1.01) 29.08 (1.57) 36.48 (2.51) 36.91 (0.36) 35.89 (0.65) 39.80 (1.42) 39.00 (0.87) Llama3.1-8B 27.16 (1.62) 25.76 (1.42) 30.61 (1.25) 29.59 (1.25) 30.91 (0.99) 29.76 (1.26) 31.59 (0.69) Mixtral-8x7B 34.19 (0.68) 25.23 (1.19) 37.72 (0.96) 29.48 (0.87) 29.09 (1.46) 31.78 (1.71) 29.45 (0.04) Mistral-7B 27.78 (1.62) 25.02 (1.26) 30.65 (1.39) 24.99 (1.51) 27.34 (0.49) 27.97 (1.31) 26.97 (1.45) Gemma-7B 17.98 (2.15) 21.64 (0.39) 22.31 (2.15) 18.66 (1.55) 25.97 (1.92) 21.79 (0.40) 21.22 (0.59) Gemma-2B 1.04 (0.19) 24.19 (0.82) 25.25 (1.02) 24.32 (1.55) 25.03 (0.66) 25.44 (1.96) 23.62 (0.36)
$\dagger$: Llama3.1 supports up to 128k tokens, but we limited it to 8k tokens due to GPU VRAM limits.
@@
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. Open Tülu 2-70B 0.21 (0.07) 18.24 (0.84) 20.60 (1.00) 18.64 (1.81) 31.71 (2.22) 7.82 (1.57) 17.31 (0.61) Tülu 2-7B 0.74 (0.15) 13.19 (0.35) 19.54 (1.29) 26.07 (2.00) 13.87 (0.71) 18.35 (1.21) 27.48 (2.04) Llama3.1-70B 0.64 (0.10) 29.29 (0.59) 0.60 (0.12) 15.07 (5.12) 39.08 (0.99) 2.43 (0.10) 18.18 (0.16) Llama3.1-8B 26.61 (1.24) 26.86 (0.78) 31.20 (1.87) 27.08 (0.63) 24.82 (1.07) 31.72 (1.66) 22.69 (0.63) Mixtral-8x7B 2.41 (0.40) 14.90 (0.82) 23.55 (0.40) 15.64 (0.47) 16.43 (1.68) 22.95 (0.68) 13.22 (1.61) Mistral-7B 3.35 (0.58) 24.44 (1.13) 31.39 (0.70) 24.26 (1.60) 29.82 (0.95) 30.21 (0.90) 23.90 (0.51) Gemma-7B 18.05 (0.97) 22.52 (0.81) 20.64 (0.26) 16.63 (1.59) 23.41 (1.26) 18.34 (0.82) 19.48 (2.45) Gemma-2B 0.47 (0.13) 24.31 (0.96) 24.77 (0.74) 25.74 (1.46) 28.41 (1.20) 24.68 (1.45) 24.75 (1.50)
$\dagger$: Llama3.1 supports up to 128k tokens, but we limited it to 8k tokens due to GPU VRAM limits.

### v5

@@ Appendix F Prompt for Response Generation
The prompt given to the conversation system to answer questions using dialogue history is shown in Table 9. An example where the placeholders from Table 9 are filled with actual values can be found in Table 10.
@@ G.1 Time Limit
In DialSim, the time limit is a controllable parameter, giving developers the flexibility to conduct experiments with any chosen time constraint, or even without one. When a time limit is set, the experimental environment can impact performance. Consequently, depending on the environment in which the conversation system is deployed, this could serve as a criterion for selecting the system with relatively better performance. It is important to note that the primary objective of DialSim is not to evaluate the inference speed of LLMs, but rather to assess the end-to-end performance of conversation systems, where techniques like model sharding and tensor parallelism can be a part of the conversation system to decrease the response latency if needed.
@@ G.1.2 Average Time Interval Between Utterances
To account for this, we set a 6-second window as the response time limit for systems and conducted experiments to determine whether current models could meet this criterion. It is important to emphasize that the primary goal of these experiments was not to evaluate the absolute performance of the models but to showcase the range of analyses possible under time limits.
@@ G.1.3 Internet Speed
The performance of API-based models can be affected by internet speed. To analyze this, we conducted a comparative analysis of the response times between API-based models and open-source models. In our analysis of systems using OpenAI Embedding-Session Sum., we found that the API-based systems achieved average response times of 1.50 seconds for GPT-4o-mini, 1.73 seconds for GPT-3.5 and 2.69 seconds for Gemini 1.0 pro. In comparison, systems using open-source models showed average response times ranging from 2.06 seconds (Gemma 2B) to 7.15 seconds (Tulu2 70B). These results suggest that, even when accounting for both internet communication and model inference, remote API-based models are generally faster than open-source alternatives. This indicates that internet latency has a minimal impact on our evaluation.
@@ G.2 Question Format
LongDialQA is a dataset that includes pairs of questions, answers, and choices. The questions are available in three formats: template-based multiple-choice, natural language multiple-choice, and open-ended. Users can choose any of these formats to evaluate the system’s performance.
@@
Additionally, we offer the option to ask questions in an open-ended format (e.g., “Who was going out with Paul in September 1994?”) without providing answer choices. This approach allows us to evaluate the system’s ability to generate open-ended responses. The open-ended format is particularly useful for fan quiz-based questions, where some answers may require longer responses (e.g., Question: “Why did Monica and Chandler say they were late getting to the hospital?” Correct answer: “Monica went back for her jacket”).
@@ G.3 Number of Retrieved Dialogue History
By default, systems retrieved up to 20 utterances, 10 entire sessions, and 15 session summaries, depending on the storing method, though some LLMs with shorter context lengths retrieved fewer histories accordingly.
@@ Appendix I Experimental Results in an Unlimited Time Setting
The experimental results for different time limits are shown in Figure 6 and Figure 7. Figure 6 illustrates the performance over different time limits in the BM25-Session Entire setting, while Figure 7 displays the performance in the Oracle setting. Due to the high costs, time-based experiments with GPT-4o, Gemini-1.5 Pro, and Claude-3 Opus were conducted exclusively in the Oracle setting. One key observation from the results is the performance of GPT-3.5, GPT-4o-mini, and GPT-4o. These models demonstrated consistent performance with quick inference times, handling up to a 3-second limit in the BM25-Session Entire setting and up to a 1-second limit in the Oracle setting. Consequently, these models are optimal for tasks requiring real-time communication without delays.
@@
Figure 4: The result of asking GPT-4o to explain Season 2, Episode 7 of Friends.
Figure 5: The result of asking GPT-4o to explain Season 3, Episode 14 of Friends.
@@
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 22.68 (2.12) 19.77 (2.02) 36.63 (1.82) 30.10 (1.44) 29.54 (0.71) 32.34 (0.58) 35.72 (0.81) GPT-3.5 32.49 (1.72) 25.32 (1.20) 35.59 (2.12) 33.86 (1.09) 27.81 (0.40) 32.97 (0.86) 37.02 (1.13) Gemini 1.0 pro 3.49 (0.69) 25.87 (1.23) 30.72 (0.18) 38.16 (1.25) 37.42 (0.68) 32.09 (0.44) 36.30 (0.32) Open Tülu 2-70B 0.62 (0.13) 21.08 (0.70) 18.95 (1.07) 22.36 (0.65) 34.64 (0.69) 9.08 (1.00) 20.22 (1.48) Tülu 2-7B 0.53 (0.18) 15.58 (1.34) 22.26 (0.53) 29.99 (0.57) 16.84 (2.13) 21.48 (0.77) 28.69 (1.15) Llama3.1-70B 0.25 (0.07) 21.55 (0.93) 0.15 (0.12) 1.26 (0.31) 34.21 (1.59) 3.89 (1.05) 14.89 (1.74) Llama3.1-8B 21.30 (1.68) 12.80 (1.06) 25.50 (0.16) 18.56 (0.99) 23.10 (2.69) 25.48 (3.65) 20.75 (1.58) Mixtral-8x7B 1.95 (0.34) 15.91 (0.71) 34.52 (1.12) 16.83 (1.60) 17.45 (0.49) 34.98 (0.99) 13.83 (2.18) Mistral-7B 3.11 (0.21) 24.69 (1.82) 34.26 (0.60) 32.17 (1.39) 30.23 (0.62) 33.36 (0.56) 29.19 (1.54) Gemma-7B 16.40 (0.74) 21.40 (2.33) 19.74 (2.45) 16.67 (0.40) 24.50 (1.87) 20.22 (1.39) 16.12 (0.52) Gemma-2B 1.56 (0.06) 28.94 (0.35) 26.12 (2.22) 33.47 (1.41) 27.92 (0.68) 29.40 (1.79) 34.86 (3.20)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
@@
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 28.48 (1.01) 29.44 (0.62) 43.16 (1.37) 35.92 (2.50) 37.81 (0.30) 40.91 (0.37) 42.83 (1.12) GPT-3.5 36.54 (0.32) 36.63 (0.57) 45.33 (1.00) 40.93 (0.13) 42.49 (1.24) 43.04 (0.82) 45.18 (0.56) Gemini 1.0 pro 2.42 (0.18) 35.11 (0.50) 48.90 (1.57) 40.91 (0.75) 44.72 (0.19) 46.63 (0.89) 45.82 (0.97) Open Tülu 2-70B 0.46 (0.09) 22.33 (1.00) 35.52 (0.89) 23.49 (1.16) 38.61 (1.02) 43.49 (1.27) 23.54 (0.52) Tülu 2-7B 0.32 (0.04) 25.86 (0.54) 27.95 (1.03) 36.60 (2.11) 22.13 (0.33) 29.50 (0.56) 35.51 (1.40) Llama3.1-70B 0.19 (0.07) 29.21 (0.56) 13.31 (0.94) 21.32 (5.22) 47.41 (0.93) 47.07 (1.32) 19.46 (1.73) Llama3.1-8B 21.87 (0.60) 22.03 (0.32) 37.94 (1.28) 29.16 (1.80) 27.76 (3.52) 37.67 (1.70) 26.67 (0.83) Mixtral-8x7B 1.53 (0.41) 19.63 (0.79) 34.35 (1.19) 16.07 (0.56) 20.02 (0.44) 30.44 (1.69) 12.43 (1.04) Mistral-7B 2.55 (0.09) 30.65 (0.45) 41.16 (1.26) 35.67 (1.68) 36.92 (2.13) 42.71 (1.24) 37.65 (2.42) Gemma-7B 17.81 (0.86) 21.58 (0.61) 25.62 (0.02) 12.20 (0.57) 24.88 (0.93) 24.38 (0.52) 15.70 (0.43) Gemma-2B 0.83 (0.16) 29.71 (0.69) 28.11 (1.14) 34.63 (0.94) 31.54 (0.65) 30.31 (0.16) 33.37 (0.27)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
@@
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 38.91 (0.99) 34.44 (0.52) 49.21 (0.12) 42.23 (1.57) 38.91 (0.74) 43.64 (0.42) 42.40 (0.99) GPT-3.5 31.81 (1.33) 26.91 (2.30) 39.45 (1.40) 32.77 (1.31) 32.41 (0.96) 35.78 (0.74) 35.98 (1.75) Gemini 1.0 pro 28.36 (0.97) 28.10 (1.08) 39.90 (1.08) 34.11 (1.64) 34.26 (2.91) 30.93 (2.17) 33.96 (2.11) Open Tülu 2-70B 3.31 (0.32) 29.87 (0.65) 35.87 (2.56) 34.72 (1.63) 37.07 (0.72) 33.63 (1.32) 38.62 (1.94) Tülu 2-7B 1.57 (0.12) 28.93 (2.81) 28.72 (1.80) 30.86 (2.29) 34.55 (0.47) 31.04 (0.96) 32.12 (0.75) Llama3.1-70B 36.36 (0.68) 31.84 (1.29) 43.17 (0.99) 43.81 (0.94) 39.85 (2.08) 43.17 (0.68) 48.49 (0.97) Llama3.1-8B 28.78 (0.34) 29.89 (1.56) 34.70 (1.75) 33.93 (1.76) 31.63 (2.17) 32.91 (0.51) 35.59 (1.09) Mixtral-8x7B 42.19 (1.76) 31.84 (0.78) 46.47 (1.75) 32.31 (1.09) 35.51 (0.19) 41.24 (2.90) 34.18 (0.96) Mistral-7B 32.93 (0.59) 28.20 (1.17) 35.09 (1.76) 30.16 (1.82) 30.12 (1.45) 31.00 (1.93) 30.80 (1.75) Gemma-7B 18.78 (0.87) 22.26 (1.52) 23.62 (2.09) 19.83 (1.74) 25.07 (0.49) 22.48 (0.25) 20.08 (0.76) Gemma-2B 1.16 (0.26) 25.03 (1.54) 24.64 (1.31) 24.84 (2.05) 28.06 (1.38) 24.56 (2.60) 28.28 (1.94)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
@@
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 31.37 (1.57) 27.54 (1.85) 40.36 (0.73) 29.71 (0.84) 33.63 (0.30) 35.93 (1.56) 37.30 (1.71) Open Tülu 2-70B 0.31 (0.13) 17.03 (0.94) 15.20 (0.87) 18.45 (1.04) 26.89 (0.54) 6.92 (0.47) 12.86 (1.23) Tülu 2-7B 0.73 (0.29) 12.50 (1.73) 17.58 (1.14) 24.21 (1.09) 10.20 (0.21) 14.26 (0.92) 21.03 (0.58) Llama3.1-70B 0.51 (0.00) 27.84 (1.89) 0.60 (0.06) 13.84 (2.31) 35.67 (1.89) 1.53 (028) 20.90 (0.37) Llama3.1-8B 25.84 (1.16) 25.24 (0.30) 28.86 (1.01) 24.56 (0.99) 28.86 (1.10) 32.35 (1.51) 24.05 (1.36) Mixtral-8x7B 1.77 (0.19) 14.03 (0.12) 21.11 (1.07) 15.50 (0.68) 13.14 (0.83) 18.03 (0.44) 18.47 (0.55) Mistral-7B 2.34 (0.17) 22.29 (1.43) 27.08 (0.99) 24.15 (1.76) 25.17 (1.74) 26.76 (2.64) 23.81 (2.53) Gemma-7B 18.87 (1.43) 22.85 (0.81) 22.96 (1.34) 17.95 (0.62) 25.46 (2.08) 21.53 (1.00) 17.66 (1.31) Gemma-2B 0.78 (0.22) 22.99 (0.66) 25.48 (1.54) 25.86 (2.48) 25.08 (1.34) 25.21 (0.22) 26.14 (1.71)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
@@
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 32.18 (1.74) 30.27 (0.73) 38.40 (0.30) 34.70 (0.22) 32.90 (0.57) 36.10 (0.36) 36.40 (0.73) Open Tülu 2-70B 2.54 (0.21) 26.47 (1.91) 31.75 (1.71) 30.94 (2.41) 31.90 (1.25) 29.83 (1.03) 31.86 (2.13) Tülu 2-7B 1.15 (0.06) 28.20 (1.63) 27.64 (2.37) 27.78 (1.32) 28.98 (0.96) 25.03 (0.94) 29.08 (2.47) Llama3.1-70B 31.38 (1.01) 29.08 (1.57) 36.48 (2.51) 36.91 (0.36) 35.89 (0.65) 39.80 (1.42) 39.00 (0.87) Llama3.1-8B 27.16 (1.62) 25.76 (1.42) 30.61 (1.25) 29.59 (1.25) 30.91 (0.99) 29.76 (1.26) 31.59 (0.69) Mixtral-8x7B 34.19 (0.68) 25.23 (1.19) 37.72 (0.96) 29.48 (0.87) 29.09 (1.46) 31.78 (1.71) 29.45 (0.04) Mistral-7B 27.78 (1.62) 25.02 (1.26) 30.65 (1.39) 24.99 (1.51) 27.34 (0.49) 27.97 (1.31) 26.97 (1.45) Gemma-7B 17.98 (2.15) 21.64 (0.39) 22.31 (2.15) 18.66 (1.55) 25.97 (1.92) 21.79 (0.40) 21.22 (0.59) Gemma-2B 1.04 (0.19) 24.19 (0.82) 25.25 (1.02) 24.32 (1.55) 25.03 (0.66) 25.44 (1.96) 23.62 (0.36)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
@@
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 31.72 (1.69) 29.29 (0.74) 42.95 (2.04) 31.12 (0.94) 35.08 (1.02) 38.14 (3.63) 36.91 (2.00) Open Tülu 2-70B 0.21 (0.07) 18.24 (0.84) 20.60 (1.00) 18.64 (1.81) 31.71 (2.22) 7.82 (1.57) 17.31 (0.61) Tülu 2-7B 0.74 (0.15) 13.19 (0.35) 19.54 (1.29) 26.07 (2.00) 13.87 (0.71) 18.35 (1.21) 27.48 (2.04) Llama3.1-70B 0.64 (0.10) 29.29 (0.59) 0.60 (0.12) 15.07 (5.12) 39.08 (0.99) 2.43 (0.10) 18.18 (0.16) Llama3.1-8B 26.61 (1.24) 26.86 (0.78) 31.20 (1.87) 27.08 (0.63) 24.82 (1.07) 31.72 (1.66) 22.69 (0.63) Mixtral-8x7B 2.41 (0.40) 14.90 (0.82) 23.55 (0.40) 15.64 (0.47) 16.43 (1.68) 22.95 (0.68) 13.22 (1.61) Mistral-7B 3.35 (0.58) 24.44 (1.13) 31.39 (0.70) 24.26 (1.60) 29.82 (0.95) 30.21 (0.90) 23.90 (0.51) Gemma-7B 18.05 (0.97) 22.52 (0.81) 20.64 (0.26) 16.63 (1.59) 23.41 (1.26) 18.34 (0.82) 19.48 (2.45) Gemma-2B 0.47 (0.13) 24.31 (0.96) 24.77 (0.74) 25.74 (1.46) 28.41 (1.20) 24.68 (1.45) 24.75 (1.50)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.